# On the Origin of the High Star-Formation Efficiency in Massive Galaxies at Cosmic Dawn

### v1

@@ Abstract
Motivated by the early excess of bright galaxies seen by JWST, we run zoom-in cosmological simulations of a massive galaxy at Cosmic Dawn (MDG), in a halo of $10^{11}M_{\odot}$ at $z=9$, using the hydro-gravitational code RAMSES at an effective resolution $\sim 10~{}{\rm pc}$. We investigate physical mechanisms that enhance the star-formation efficiencies (SFEs) under the unique conditions of high gas density ($\sim 3\times 10^{3}~{}{\rm cm^{-3}}$, $\sim 10^{4}~{}M_{\odot}/{\rm pc^{2}}$). Our fiducial star formation recipe uses a physically-motivated, turbulence-based, multi-freefall model, avoiding ad-hoc extrapolation from lower redshifts. By $z=9$, our simulated galaxy is a clumpy, thick, rotating disc with a high stellar mass of a few $10^{9}~{}M_{\odot}$ and high star formation rate of $\sim 100~{}M_{\odot}/{\rm yr}$. The high gas density makes supernova (SN) feedback less effective at suppressing star formation, producing a relatively high local SFE $\gtrsim 10\%$. The global SFE is dominated by feedback-driven outflows and is only weakly correlated with the local SFE. Photoionization heating can enhance the effects of SN feedback on the local SFE by making more SNe explode in diffuse environments, but the global SFE remains high even in our simulations with the strongest feedback. Intense accretion at Cosmic Dawn seeds strong turbulence, which reduces the local SFE for the same gas conditions due to turbulent pressure support. This prevents star-forming clouds from catastrophically collapsing, but this only weakly affects the global SFE. The star formation histories of our simulated galaxies are in the ballpark of the MDGs seen by JWST, despite our limited resolution. They set the stage for future simulations which treat radiation self-consistently and use a higher effective resolution $\sim 1~{}{\rm pc}$ which captures the physics of star-forming clouds.
@@ 1 Introduction
The James Webb Space Telescope (JWST) has opened a new frontier for empirical constraints on galaxy formation models at cosmic dawn ($z\geq 9$), an epoch starting with the formation of the first stars and ending with the reionization of the inter-galactic medium (IGM). At these redshifts, JWST’s NIRCam and NIRSpec instruments probe the rest-frame ultraviolet (UV), where they can provide estimates for stellar mass and star formation rate (SFR).
@@
So far, JWST has found a growing list of $z\sim 10$ objects with high stellar masses $M_{*}\gtrsim 10^{8}~{}M_{\odot}$ and/or high SFRs $\dot{M}_{*}\gtrsim 1~{}M_{\odot}/{\rm yr}$ (e.g. Finkelstein et al., 2023b; Labbé et al., 2023; Mason et al., 2023; Carniani et al., 2024). These massive cosmic dawn galaxies (MDGs) form an excess in the abundance of the (UV-)brightest galaxies relative to expectations of standard galaxy formation models (Boylan-Kolchin, 2023). Shen et al. (2023) show that to reproduce the UV luminosity function (UVLF) inferred from MDGs, one requires that $\sim 30\%$ of the baryonic matter accreting onto a dark matter (DM) halo is converted into stars. This is far higher than the canonical value of $\lesssim 5\%$ for galaxies at lower redshift in a similar mass and redshift range.
@@
The explanations for this tension fall into four categories. First, star formation might be intrinsically more efficient at Cosmic Dawn (Dekel et al., 2023). Second, a low mass-to-light ratio at Cosmic Dawn could cause models calibrated to the lower-redshift Universe to overestimate stellar masses. These explanations include the presence of AGN (Wang et al., 2024), a top-heavy stellar initial mass function (IMF) (Sharda & Krumholz, 2022; Cameron et al., 2023), a lack of dust along the line of sight (Mason et al., 2023; Ferrara et al., 2023), and differences in dust properties. Third, observations could be biased by bursty star formation at Cosmic Dawn (Sun et al., 2023; Shen et al., 2023). In this scenario, the bright end of the UVLF is dominated by less intrinsically bright galaxies which have undergone a recent burst. Finally, a modified $\Lambda$-CDM cosmology can produce a greater abundance of massive haloes at Cosmic Dawn. These explanations include “Early Dark Energy” (Klypin et al., 2021) or massive primordial black hole seeds (Liu & Bromm, 2022).
@@
Some combination of effects may contribute to the tension. In this work, we explore the possibility that star formation is intrinsically more efficient at Cosmic Dawn. Ultimately, a change in the physics of star formation must result from differences in the properties of the interstellar medium (ISM). A key difference is that the typical density of the ISM was greater at Cosmic Dawn by a factor $\sim 1000$ due to cosmic expansion.
@@
In Figure 1, we plot the comoving number density of dark matter haloes as a function of halo mass and redshift, assuming cosmological parameters from the Plank 2018 results (Planck Collaboration et al., 2020). We estimate the number density by integrating the Press-Schechter halo mass function (Press & Schechter, 1974) using the Python package COLOSSUS (Diemer, 2018). haloes become increasingly rare at larger masses and higher redshifts. We also plot a simple estimate for the baryon surface density of the galaxy hosted by the halo (App. A).
The MDGs seen by JWST must have number densities on the order of a few per effective JWST survey volume, which is roughly $n_{\rm halo}\sim 10^{-6}~{}(h_{0}^{-1}{\rm cMpc^{3}})^{-3}$. For example, $V_{\rm eff}(M\leq-20)\simeq 160000~{}{\rm cMpc^{3}}$ for the Cosmic Evolution Early Release Science (CEERS Finkelstein et al., 2023a, ERS-1345). Using Figure 1, this implies that at a redshift $z\sim 9$, the MDGs seen by JWST have halo masses $M_{\rm halo}\sim 10^{11}~{}M_{\rm halo}$ and baryon surface densities $\Sigma_{\rm gas}\sim 10^{4}~{}M_{\odot}/{\rm pc^{2}}$, two orders of magnitude larger than the typical surface densities in Milky Way like galaxies (Ostriker & Kim, 2022). For a disc scale height $H/R\sim 0.1$, this gives a Hydrogen number density $n_{\rm H}\sim 3\times 10^{3}~{}{\rm cm^{-3}}$. The density in star-forming regions is expected to be up to an order of magnitude higher $n_{\rm H}\sim 3\times 10^{4}~{}{\rm cm^{-3}}$.
@@
Figure 1: The comoving number density of dark matter haloes as a function of halo mass and redshift. A simple estimate for gas surface density (App. A) is used to plot lines of constant gas surface density in red. The most massive galaxies at Cosmic Dawn seen by JWST likely have number densities $n_{\rm halo}\sim 10^{-6}~{}(h_{0}^{-1}{\rm cMpc^{3}})^{-3}$. Our simulation (red star) probes galaxies with similar number densities. The high densities $\Sigma_{\rm gas}\sim 10^{4}~{}M_{\odot}/{\rm pc^{2}}$ and $n_{\rm H}\sim 3\times 10^{3}~{}{\rm cm^{-3}}$ create a unique environment for star formation which is not realized in the lower redshift Universe.
@@
Such high densities create a unique environment for star formation which is not realized in the lower redshift Universe, except in nuclear regions of lower redshift galaxies, where the typical result is a starburst (e.g. Naab & Ostriker, 2017). This suggests that MDGs might have starburst-like SFR densities throughout their entire volume, rather than just in their center. In their feedback-free starburst (FFB) model, Dekel et al. (2023) argue that the high density conditions of Cosmic Dawn might suppress the stellar feedback processes which usually regulate star formation.
@@
However, the physical conditions of the ISM in MDGs differs from present day galaxies in other ways too. The rapid accretion of cold streams in massive galaxies at high redshift is expected to seed strong turbulence (Dekel et al., 2009; Ginzburg et al., 2022). This is empirically supported by the high velocity dispersions observed in galaxies at $z\gtrsim 6$ de Graaff et al. (2024b). Turbulent pressure may support molecular clouds against gravitational collapse, suppressing star formation locally. Turbulence can also reduce the clustering of supernovae (SNe), making SN feedback less effective (Gentry et al., 2017). In addition to turbulence, the low metal content of the ISM and the higher CMB temperatur at Cosmic Dawn may suppress star formation by making cooling less efficient.
We capture these effects in a set of cosmological simulations from $z=100$ to $z=9$, focusing on a halo which reaches a mass $M_{\rm halo}=10^{11}~{}M_{\odot}$ at $z=9$, indicated by the red star in Figure 1. Our models vary in their treatment of turbulence and stellar feedback, allowing us to investigate how star formation on a global scale is connected to physics on a local scale.
The star formation recipe commonly used in simulations of galaxy formation defines the local SFR density $\dot{\rho}_{*}=\epsilon_{\rm ff}\rho/\tau_{\rm ff}$, where $\epsilon_{\rm ff}$ is the local star formation efficiency (SFE) per freefall time $\tau_{\rm ff}$. Typically, $\epsilon_{\rm ff}$ is set to a constant value on the order of a few percent, consistent with observations of the local Universe (e.g. Hopkins et al., 2014). However, to understand how the different physical conditions in the ISM at Cosmic Dawn affect star formation, we need a more predictive model. Our fiducial model uses a physically-motivated, turbulence-based, multi-freefall model (Federrath & Klessen, 2012), avoiding ad hoc extrapolation from lower redshifts.
@@
This work is only the first step towards understanding the physics of star formation in MDGs. Our treatment of stellar feedback is crude due to the limitations of our effective resolution $\sim 10~{}{\rm pc}$ in a pure hydrodynamics simulation (i.e. no radiation). Therefore, our quantitative results should be taken with a grain of salt. Future work will treat radiation self-consistently and use a higher effective resolution $\sim 1~{}{\rm pc}$ which captures the physics of star-forming clouds.
@@
In Section 2, we explain our numerical methods and our suite of models. In Section 3, we analyze our results. In Section 4, we interpret our results with toy models, discuss caveats, and discuss observational implications. We conclude in Section 5. Tangential discussions are provided in the appendices.
@@ 2.1 RAMSES
Each particle is defined by a unique identification number (ID), a position $\mathbf{x}$ and a velocity $\mathbf{v}$. Star cluster particles carry additional metadata including their mass $m_{*}$, metallicity $Z$, and time of birth $t_{\rm birth}$. We evolve density $\rho$, pressure $P$, velocity $\mathbf{v}$, and metallicity $Z$ on the grid according to the Euler equations using the MUSCL-Hancock scheme, a second-order Godunov method, and the Harten-Lax-van Leer-Contact Leer (HLLC) Riemann solver (Harten et al., 1983). We use a fine timestep controlled by several stability conditions (see Teyssier, 2002) and whose size is roughly $\tau_{\rm fine}\simeq 500~{}{\rm yr}$.
We assume an ideal gas equation of state appropriate for a rarefied plasma with $\gamma=5/3$. The turbulent kinetic energy (TKE) $\varepsilon_{\rm turb}$ and a refinement mask $M$ are advected passively with the flow. The trajectories of dark matter and star cluster particles are computed using a particle-mesh solver. The Poisson equation is solved using a multi-grid method (Guillet & Teyssier, 2011) with Dirichlet boundary conditions at level boundaries.
The grid resolution can be defined in terms of a refinement level $\ell$ as $\Delta x=L/2^{\ell}$. Similarly, we can define the mass resolution of dark matter particles as
@@
We generate initial conditions for a low-resolution ($\ell=9$) dark matter only simulation in a comoving volume $L^{3}=(100h_{0}^{-1}{\rm cMpc})^{3}$ at $z=100$ using MUSIC (Hahn & Abel, 2011). We run the simulation until $z=9$ and identify a candidate halo with a mass $M_{\rm halo}\simeq 10^{11}~{}M_{\odot}$. We trace the dark matter particles of the candidate halo back to their positions at $z=100$ and define the convex hull created by these particles as the zoom region.
@@
Criterion (i) ensures that only cells which exchange gas with the initial zoom-in region are refined. Criterion (iv) ensures that before a cell acquires enough mass to become gravitationally unstable at a given resolution, the cell is refined. Therefore, only cells at the highest refinement level should become gravitationally unstable and form stars.
@@
We do not directly model Pop III star formation. Instead, we enrich the ISM to $10^{-3}Z_{\odot}$ in the initial conditions. Previous simulation work shows that a single pair instability SN can enrich the host halo to at least this metallicity (Wise et al., 2011). Furthermore, this is above the critical metallicity $\sim 10^{-3.5}Z_{\odot}$ at which Pop III stars cease to form due to fine-structure Carbon and Oxygen line cooling (Bromm & Loeb, 2003). Therefore, our initial conditions are consistent with our lack of Pop III star modeling.

### v2

@@ Abstract
Motivated by the early excess of bright galaxies seen by JWST, we run zoom-in cosmological simulations of a massive galaxy at Cosmic Dawn, in a halo of $10^{11}M_{\odot}$ at $z=9$, using the hydro-gravitational code RAMSES at an effective resolution $\sim 10~{}{\rm pc}$. We investigate physical mechanisms that enhance the star-formation efficiencies (SFEs) at the high gas densities of the star-forming regions in this galaxy ($\sim 3\times 10^{3}~{}{\rm cm^{-3}}$, $\sim 10^{4}~{}M_{\odot}/{\rm pc^{2}}$). Our fiducial star formation recipe uses a physically-motivated, turbulence-based, multi-freefall model, avoiding ad hoc extrapolation from lower redshifts. By $z=9$, our simulated galaxy is a clumpy, thick, rotating disc with a high stellar mass $\sim 3\times 10^{9}~{}M_{\odot}$ and high star formation rate $\sim 50~{}M_{\odot}/{\rm yr}$. The high gas density makes supernova (SN) feedback less efficient, producing a high local SFE $\gtrsim 10\%$. The global SFE is set by feedback-driven outflows and only weakly correlated with the local SFE. Photoionization heating makes SN feedback more efficient, but the integrated SFE always remains high. Intense accretion at Cosmic Dawn seeds turbulence which reduces local SFE, but this only weakly affects the global SFE. The star formation histories of our simulated galaxies are similar to observed massive galaxies at Cosmic Dawn, despite our limited resolution. We set the stage for future simulations which treat radiation self-consistently and use a higher effective resolution $\sim 1~{}{\rm pc}$ that captures the physics of star-forming clouds.
@@ 1 Introduction
The James Webb Space Telescope (JWST) has opened a new frontier for empirical constraints on galaxy formation models at Cosmic Dawn ($z\geq 9$), an epoch starting with the formation of the first stars and ending with the reionization of the inter-galactic medium (IGM). At these redshifts, JWST’s NIRCam and NIRSpec instruments probe the rest-frame ultraviolet (UV), where they provide estimates for stellar mass and star formation rate (SFR).
@@
So far, JWST has found a growing list of galaxies at Cosmic Dawn with high stellar masses $M_{*}\gtrsim 10^{8}~{}M_{\odot}$ and/or high SFRs $\dot{M}_{*}\gtrsim 1~{}M_{\odot}/{\rm yr}$ (e.g. Finkelstein et al., 2023; Labbé et al., 2023; Mason et al., 2023; Carniani et al., 2024). These massive Dawn galaxies (MDGs) constitute an excess in the abundance of the (UV-)brightest galaxies relative to expectations of standard galaxy formation models (Boylan-Kolchin, 2023). Shen et al. (2023) show that to reproduce the UV luminosity function (UVLF) inferred from MDGs, one requires that $\sim 30\%$ of the baryonic matter accreting onto a dark matter (DM) halo is converted into stars. This is far higher than the canonical value of $\lesssim 5\%$ inferred for galaxies at lower redshift based on abundance matching of stellar masses and $\Lambda{\rm CDM}$ haloes (Rodríguez-Puebla et al., 2017; Moster et al., 2018; Behroozi et al., 2019).
@@
The explanations for this tension fall into four categories. First, star formation might be intrinsically more efficient at Cosmic Dawn (Dekel et al., 2023). Second, a low mass-to-light ratio at Cosmic Dawn might cause models calibrated to the lower-redshift Universe to overestimate stellar masses. These explanations include the presence of active galactic nuclei (AGN) (Wang et al., 2024), a top-heavy stellar initial mass function (IMF) (Sharda & Krumholz, 2022; Cameron et al., 2023), a lack of dust along the line of sight (Mason et al., 2023; Ferrara et al., 2023), and differences in dust properties. Third, observations could be biased by bursty star formation at Cosmic Dawn (Sun et al., 2023; Shen et al., 2023). In this scenario, the bright end of the UVLF is dominated by less intrinsically bright galaxies which have undergone a recent burst. Finally, a modified $\Lambda$-CDM cosmology can produce a greater abundance of massive haloes at Cosmic Dawn. These explanations include “Early Dark Energy” (Klypin et al., 2021; Shen et al., 2024) or massive primordial black hole seeds (Liu & Bromm, 2022).
@@
Some combination of effects may contribute to the tension. In this work, we explore the possibility that star formation is intrinsically more efficient at Cosmic Dawn due to differences in the properties of the interstellar medium (ISM). The mean baryon density of the Universe scales as $(1+z)^{-3}$ due to cosmic expansion, suggesting that massive dark matter haloes at Cosmic Dawn may host galaxies with high gas densities.
@@
At high redshift, massive haloes become increasingly rare. Typical comoving volumes of JWST galaxy surveys are $\sim 10^{5}$ to $10^{6}~{}{\rm cMpc^{3}}$. For example, in the Cosmic Evolution Early Release Science (CEERS Finkelstein et al., 2024, ERS-1345), the effective survey volume at UV magnitude $\leq-20$ is $\simeq 160000~{}{\rm cMpc^{3}}$. Thus, the haloes of observed galaxies should have comoving number densities $\gtrsim 10^{-6}~{}(h_{0}^{-1}{\rm cMpc})^{-3}$. In Figure 1, we plot the comoving number density of dark matter haloes as a function of halo mass and redshift, assuming cosmological parameters from the Plank 2018 results (Planck Collaboration et al., 2020). We estimate the number density by integrating the Press-Schechter halo mass function (Press & Schechter, 1974) using the Python package COLOSSUS (Diemer, 2018). We also plot an estimate for the typical baryon surface density of the galaxy hosted by the halo. The calculation is described in Appendix A.
Haloes at redshifts $z\sim 9$ and masses $\sim 10^{11}~{}M_{\odot}$ (marked by the red star) are interesting objects because they have number densities $\gtrsim 10^{-6}~{}(h_{0}^{-1}{\rm cMpc})^{-3}$ and they are expected to host galaxies with high baryon surface densities $\sim 10^{4}~{}M_{\odot}/{\rm pc^{2}}$. For comparison, a Milky-Way-like galaxy at $z=0$ with halo mass $10^{12}~{}M_{\odot}$ has the baryon surface density $\sim 3\times 10^{2}~{}M_{\odot}/{\rm pc^{2}}$. We expect the difference in surface densities to be even greater for the gas component, because at low redshift, massive galaxies tend to be quenched with most of their baryons in stars. We can also estimate the typical Hydrogen number density by assuming a disc scale height (App. A). For a scale height $(H/R)_{\rm gal}=0.1$ and 50% of the baryon mass in gas, the number density is $\sim 3\times 10^{3}~{}{\rm cm^{-3}}$.
@@
Figure 1: The comoving number density of dark matter haloes as a function of halo mass and redshift. A simple estimate for baryon surface density (App. A) is used to plot lines of constant baryon surface density in red. The most massive galaxies at Cosmic Dawn seen by JWST likely have comoving number densities $\sim 10^{-6}~{}(h_{0}^{-1}{\rm cMpc^{3}})^{-3}$. Our simulation (red star) probes galaxies with similar number densities. The high baryon surface densities $\sim 10^{4}~{}M_{\odot}/{\rm pc^{2}}$ and number densities $\sim 3\times 10^{3}~{}{\rm cm^{-3}}$ create a unique environment for star formation which is not realized in the lower redshift Universe.
@@
Such high densities create a unique environment for star formation which is not realized in the lower redshift Universe, except in nuclear regions of some galaxies, where the typical result is a starburst (e.g. Naab & Ostriker, 2017). MDGs may have starburst-like SFR densities throughout an extended region, rather than just in their center. In their feedback-free starburst (FFB) model, Dekel et al. (2023) argue that the high density conditions of Cosmic Dawn suppress the stellar feedback processes which usually regulate star formation.
@@
The physical conditions in MDGs differ from present day galaxies in other ways too. The rapid accretion of cold streams in massive galaxies at high redshift is expected to seed strong turbulence (Dekel et al., 2009; Ginzburg et al., 2022). This is empirically supported by the high velocity dispersions observed in galaxies at $z\gtrsim 6$ (de Graaff et al., 2024b). Turbulent pressure supports molecular clouds against gravitational collapse, suppressing local star formation. Turbulence also reduces the clustering of supernovae (SNe), making SN feedback less effective (Gentry et al., 2017). In addition to turbulence, the low metal content of the ISM and the higher CMB temperature at Cosmic Dawn may suppress star formation by making cooling less efficient.
Galaxy formation during the epoch of reionization is often studied with big box cosmological simulations such as MassiveBlack (Di Matteo et al., 2012), CROC (Gnedin, 2014), Eagle (Schaye et al., 2015; Crain et al., 2015), BlueTides (Feng et al., 2016), Aurora (Pawlik et al., 2017), Illustris-Tng (Naiman et al., 2018; Nelson et al., 2018; Marinacci et al., 2018; Springel et al., 2018; Pillepich et al., 2018), Sphinx (Rosdahl et al., 2018), SIMBA (Davé et al., 2019), Cosmic Dawn II (Ocvirk et al., 2020), and Thesan (Kannan et al., 2022). However, it is challenging to simultaneously resolve the structure of the ISM at scales $\sim 10~{}{\rm pc}$ and use a large enough volume $\simeq(100h^{-1}{\rm cMpc})^{3}$ to sample rare massive haloes.
Previous work has gotten around this by using computationally cheap dark-matter-only simulations to select massive haloes as sources for zoom-in simulations with full hydrodynamics (Katz & White, 1993; Tormen et al., 1997), including Renaissance (O’Shea et al., 2015), FirstLight (Ceverino et al., 2017), Fire-2 (Ma et al., 2018), Flares (Lovell et al., 2021; Vijayan et al., 2021), and serra (Pallottini et al., 2022). These works generally find that massive galaxies at high redshift are associated with efficient and bursty star formation. Initial work focused on forward modeling of observations and population-level properties like the UVLF rather than the details of star formation in individual galaxies. However, recent work has investigated the factors that contribute to efficient star formation in detail. Bassini et al. (2023) ran Fire-2 simulations and found that high gas volume and surface densities naturally led to efficient star formation. Similarly, Ceverino et al. (2024) analyzed FirstLight simulations and found that star formation efficiency (SFE) is correlated with redshift and halo mass, reaching up to $\sim 10\%$ for $10^{11}~{}M_{\odot}$ haloes at $z=9$ (their Fig. 7).
In this work, we also use cosmological zoom-in simulations with the principal aim of exploring the physics of star formation and feedback in the most extreme MDGs. We simulate a halo which reaches $10^{11}~{}M_{\odot}$ at $z=9$. We vary our treatment of turbulence and stellar feedback, allowing us to connect star formation on a global scale with ISM physics on a local scale. Our simulations differs from previous works in their treatment of star formation. Most high-redshift galaxy formation simulations use a star formation recipe based on the local Kennicutt-Schmidt Law (Kennicutt, 1998), where the local SFR density is $\epsilon_{\rm ff}\rho/\tau_{\rm ff}$ and $\epsilon_{\rm ff}$ is the local star formation efficiency (SFE) per freefall time $\tau_{\rm ff}$. $\epsilon_{\rm ff}$ is typically set to a constant value on the order of a few percent, consistent with observations of the local Universe (e.g. Hopkins et al., 2014). In contrast, we use a multi-freefall model (Federrath & Klessen, 2012), a physically-motivated model which predicts the SFR using a model of subgrid turbulence. This avoids ad hoc extrapolation from lower redshifts.
@@
This work is only the first step towards understanding the physics of star formation in MDGs. Our treatment of stellar feedback is crude due to the limitations of our effective resolution $\sim 10~{}{\rm pc}$ in a pure hydrodynamics simulation. Therefore, our quantitative results should be taken with a grain of salt. Future work will treat radiation self-consistently and use a higher effective resolution $\sim 1~{}{\rm pc}$ that captures the physics of star-forming clouds.
@@
In Section 2, we explain our numerical methods and our suite of models. In Section 3, we analyze our results. In Section 4, we interpret our results with toy models and discuss caveats and observational implications. We conclude in Section 5. Tangential discussions are provided in the appendices.
@@ 2.1 RAMSES
Each particle is defined by a unique identification number (ID), a position $\mathbf{x}$ and a velocity $\mathbf{v}$. Star cluster particles carry additional metadata including their mass $m_{*}$, metallicity $Z$, and time of birth $t_{\rm birth}$. We evolve density $\rho$, pressure $P$, velocity $\mathbf{v}$, and metallicity $Z$ on the grid according to the Euler equations using the MUSCL-Hancock scheme, a second-order Godunov method, and the Harten-Lax-van Leer-Contact Leer (HLLC) Riemann solver (Harten et al., 1983). We use a fine timestep controlled by several stability conditions (see Teyssier, 2002) and whose size is roughly $\simeq 500~{}{\rm yr}$.
We assume an ideal gas equation of state (EOS) appropriate for a rarefied plasma with $\gamma=5/3$. The turbulent kinetic energy (TKE) $\varepsilon_{\rm turb}$ and a refinement mask $M$ are advected passively with the flow. The trajectories of dark matter and star cluster particles are computed using a particle-mesh solver. The Poisson equation is solved using a multi-grid method (Guillet & Teyssier, 2011) with Dirichlet boundary conditions at refinement level boundaries.
The grid resolution can be defined in terms of a refinement level $\ell$ as $\Delta x=L/2^{\ell}$, where $L$ is the size of the grid. Similarly, we can define the mass resolution of dark matter particles as
@@
We generate initial conditions for a low-resolution ($\ell=9$) dark matter only simulation in a comoving volume $(100h_{0}^{-1}{\rm cMpc})^{3}$ at $z=100$ using MUSIC (Hahn & Abel, 2011). We run the simulation until $z=9$ and identify a candidate halo with a mass $\simeq 10^{11}~{}M_{\odot}$. We trace the dark matter particles of the candidate halo back to their positions at $z=100$ and define the convex hull created by these particles as the zoom region.
@@
Criterion (i) ensures that only cells which exchange gas with the initial zoom-in region are refined. Criterion (iv) ensures that we satisfy the Truelove et al. (1997) criterion in cells that are not maximally refined, preventing artificial fragmentation. Cells that are maximally refined and Jeans unstable will rapidly form stars before they have time to fragment. Therefore, we do not require an artificial pressure floor as used in previous works (Robertson & Kravtsov, 2008; Agertz et al., 2009).
@@
We do not directly model Pop III star formation. Instead, we enrich the ISM to $10^{-3}Z_{\odot}$ in the initial conditions. Previous simulation work shows that a single pair instability SN can enrich the host halo to at least this metallicity (Wise et al., 2011). Furthermore, this is above the critical metallicity $\sim 10^{-3.5}Z_{\odot}$ at which Pop III stars cease to form due to fine-structure Carbon and Oxygen line cooling (Bromm & Loeb, 2003), so our initial conditions are consistent with no Pop III stars.